OpenStudy (anonymous):

A floor plan is given for the first floor of a new house. One inch represents 12 feet. If the walk-in closet is 1/4 in. by 1/3 in. on the floor plan, what are the actual dimensions? (Round to the nearest hundredth)

OpenStudy (anonymous):

Since 1in=12 ft we can find the dimensions like this \(\large(\frac{1}{4})12~x~(\frac{1}{3})12\)

OpenStudy (anonymous):

Can you calculate those values?

OpenStudy (tkhunny):

Multiply by 1. \(\dfrac{1}{4}\;in \cdot \dfrac{12\;ft}{1\;in}\;=\;\)?? Units are wonderful things. They WILL save you.

OpenStudy (anonymous):

Yes they are, I was just being lazy. Sorry!

OpenStudy (tkhunny):

No worries. Being formal and fancy doesn't help, either, if we don't hear from the OP.

OpenStudy (anonymous):

I make such silly mistakes too. Hope they respond soon though, I need to get some sleep

OpenStudy (anonymous):

ive got help thank you. sorry.

OpenStudy (anonymous):

It's alright

OpenStudy (tkhunny):

So sad, that you will not share your victories with us. :-)

OpenStudy (anonymous):

You say sad, yet put a smiley face. :-)

OpenStudy (tkhunny):

I'm crying inside.

OpenStudy (anonymous):

Yet it can not be seen. I wish I could do that